What is the area of this triangle?
The Bottom: 14 m
Line going vertical down triangle: 6 m

Answer by checkley75(3666)

AREA=B/2*H B=BASE (BOTTOM) & H=HEIGHT (VERTICLE DIMENTION).
AREA=14/2*6
AREA=7*6=42m^2 ANSWER FOR THIS TRIANGLE.